OpenStudy (anonymous):

solve the system of equations by graphing then classify the system 3x + y = 22 4x + 5y = 22 what is the solution set? select the correct choice below and if necessary fill in the answer box to complete your choice a. the solution is ----- type an ordered pair b. there are infinitely many solutions c. there is no solution Is the system consistent or inconsistent? Is the system dependent or independent?

OpenStudy (anonymous):

by graphing you arrange it that y= something so 1st : y=22-3x 2nd: y=(22-4x)/5 now the easiest way of graphing a line is to get 2 points and connect them. 1st x=0 y=22 and x=1 y=19 are 2 points 2nd x=0 y=22/5 and x=1 y=18/5

OpenStudy (anonymous):

if they are the same line, than there are infinitely many solutions. If they are parallel than there are no solutions. If they have an intersection than they have a solution (an ordered pair)

OpenStudy (anonymous):

Hope you can do it now and it helped.

OpenStudy (anonymous):

could you specify the points i need to graph please

OpenStudy (anonymous):

(0,22) (1, 19) and (0,22/5) (1, 18/5) first is always the x coordinate, 2nd is the y

OpenStudy (anonymous):

Thanks Andras that really helped me out

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Andras i have to pick from 4 graphs that are already connected with the points could you send a pic of one showing how u graph the points so I can compare with what I have to pick from please